OpenStudy (anonymous):

Use derrivatives to determine the intervals on which f(x)=x^3+3x^2-9x+3 is increasing and the intervals where it is decreasing.

OpenStudy (campbell_st):

you would need to find the stationary points... which needs 1st derivative then solving the 1st derivative. find the second derivative and test the stationary points solve the 2nd derivative for x.. check to see if there is a change in concavity either side of your solution... if so the solution is a point of inflection.

hartnn (hartnn):

The intervals on which dy/dx >0 , f(x) will be increasing and The intervals on which dy/dx <0 , f(x) will be decreasing . what you got for dy/dx ?

OpenStudy (anonymous):

So far After taking the first derrivative and setting it equal to zero I got 1/2 and -3/2..but that doesn't really look right.

hartnn (hartnn):

setting first derivative =0 will give you extrema, which you don't exactly want. 3x^2+6x-9 >0 (or x^2 +2x-3>0) will give you intervals on which f(x) is increasing. similarly, 3x^2+6x-9 <0 (or x^2 +2x-3<0) will give you intervals on which f(x) is decreasing.

OpenStudy (anonymous):

Okay so for the increasing one can I just select a number greater than 0 such as 9 and add three then set x^2 equal to 12 and then 2x=9?

hartnn (hartnn):

how come x^2=12 ?? x^2 +2x-3>0 (x+3)(x-1) >0 can you find intervals from this ?

hartnn (hartnn):

no ?

OpenStudy (anonymous):

Alright so I get x=-3 and x=1

hartnn (hartnn):

thats for (x+3)(x-1) = 0 and not (x+3)(x-1) >0 , you know how to solve this inequality ?

OpenStudy (anonymous):

Ohh, no I do not.

hartnn (hartnn):

for (x+3)(x-1) to be positive (>0) either both (x+3) and (x-1) must be positive or both (x+3)and (x-1) must be negative, right ?

OpenStudy (anonymous):

right.

hartnn (hartnn):

so, x+3 >0 and x-1 >0 OR x+3<0 and x-1 <0 x>-3 and x>1 OR x<-3 and x<1 got this ?

OpenStudy (anonymous):

Yess, I do :)

hartnn (hartnn):

can you simplify this : x>-3 and x>1 ?

OpenStudy (anonymous):

I don't know..

hartnn (hartnn):

okk.. x>-3 means all values of x greater than -3 x>1 means all values of x greater than 1 ' and ' means common elements so, all elements for x>1 are common only. so, x>-3 and x>1 -----> x>1 only got this ?

OpenStudy (anonymous):

Okaayy

hartnn (hartnn):

what about x<-3 and x<1 ?

OpenStudy (anonymous):

x<-3?

hartnn (hartnn):

thats correct :) so, now you have x>1 or x<-3 thats your interval for which f(x) is increasing. in interval notation, it'll be (-infinity, -3)U(1,infinity) do similar process for decreasing interval.

OpenStudy (anonymous):

Ahh...Okayy! Thank you very very much!

hartnn (hartnn):

welcome ^_^ you can ask me if you get stuck ....or want to verify your answer.